OpenStudy (tumblewolf):

I need help with basic limits Our class was on snow break today and we have homework due tomorrow and we haven't covered the material. the problem is Find the M(tan) of F(x)=2x^2-x at x=3 give the equation of m(tan) I know the formula is M(tan) + lim as H approaches 0(f(a+h)-f(a)/h) but I don't know how to plug things in. Thank you!

OpenStudy (tumblewolf):

@DanJS can you please help?

OpenStudy (danjs):

ok

OpenStudy (tumblewolf):

I kinda get the concept, I just have no idea where everything goes

OpenStudy (danjs):

you need to get the derivative from the limit definition for that function?

OpenStudy (tumblewolf):

It's wanting the equation

OpenStudy (danjs):

\[F'(x) = \lim_{h \rightarrow 0}\frac{ F(x+h) - F(x) }{ h }\]

OpenStudy (danjs):

F(x) = 2x^2 - x \[F'(x) = \lim_{h \rightarrow 0}\frac{ [2*(x+h)^2 - (x+h)] - [2x^2 - x] }{ h }\]

OpenStudy (danjs):

just put in the F(x+h) and the F(x) into that, now expand it all out...

OpenStudy (danjs):

\[\lim_{h \rightarrow 0}\frac{ 2*(x^2+2xh+h^2) -x - h - 2x^2 - x }{ h }\]

OpenStudy (danjs):

\[\lim_{h \rightarrow 0}\frac{ 2x^2+4xh+2h^2 -x - h - 2x^2 - x }{ h }\]

OpenStudy (danjs):

sorry the last x in the numerator should be +x simplifies more to \[\lim_{h \rightarrow 0}\frac{ 4xh+2h^2 - h }{ h }\]

OpenStudy (danjs):

factor the h out in the top and it will cancel out with the bottom \[\lim_{h \rightarrow 0}\frac{ h*(4x+2h - 1) }{ h }\]

OpenStudy (danjs):

\[\lim_{h \rightarrow 0} (4x + 2h - 1) = 4x - 1\] F ' (x) = 4x - 1

OpenStudy (tumblewolf):

So do I not need the three?

OpenStudy (danjs):

all that was using the limit definition of the derivative to get the derivative function... F(x) = 2x^2 - x F ' (x) = 4x - 1 Is M(tan) meaning, the slope of the tangent line at x=3?, never seen the "M(tan)"

OpenStudy (tumblewolf):

Yes, that was just the easiest way to define it. Would I just use slope form?

OpenStudy (danjs):

the derivative function is the slope of a tangent line to F(x) at any given point on the function. The instantaneous rate of change of F at any value of x.

OpenStudy (tumblewolf):

So I plug into the 4(3)-1?

OpenStudy (danjs):

so if x=3 F ' (x) = 4*3 - 1 = 11 The slope of a tangent line to F(x) = 2x^2 - x at he value x=3, F ' (3) = 11

OpenStudy (tumblewolf):

I've never had to do this before so it's just like a massive overkill. So the equation would be 4(3)-1?

OpenStudy (danjs):

The derivative of F(x) = 2x^2 - x is F ' (x) = 4x - 1 that is the slope at any x value on the curve F(x). When x is 3, the slope of the tangent line there is F ' (3)..

OpenStudy (danjs):

oh, the equation for the tangent line , i see line with slope F ' (3) = 11, and through point (3 , F(3)) = (3 , 15)

OpenStudy (danjs):

y - y1 = m*(x - x1) y - 15 = 11*(x - 3) tangent line

OpenStudy (tumblewolf):

so y-11x-48?

OpenStudy (tumblewolf):

y=11x-48

OpenStudy (danjs):

if you want it in that form , y - 15 = 11x - 33 y = 11x - 18

OpenStudy (danjs):

just added the wrong way

OpenStudy (tumblewolf):

Will my basic formula always be the 4xh+2h^2-h/h ?

OpenStudy (danjs):

it depends on the function you have... The definition is this.. Slope of the tangent line at a point on the curve f(x) -- \[m = F'(x) = \lim_{h \rightarrow 0}\frac{ F(x+h) - F(x) }{ h }\] Similar to the rise/run idea from algebra, except here we have the limit as the run (change in x value, h) approaches zero. And you end up with the slope at that point tangent to the graph.

OpenStudy (danjs):

This is like the foundation idea for differential calc